OpenStudy (anonymous):

-24u^4v^3x^5 ------------------- -18uv^5x^2 simplify

OpenStudy (anonymous):

could someone please help simplify this equation?

OpenStudy (anonymous):

\[\frac{-24u^4v^3x^5}{-18uv^5x^2}=\frac {4u^3x^3}{3v^2}\rightarrow \frac {4(ux)^3}{3v^2}\]